OpenStudy (anonymous):

Evaluate i^12 I want to know how to work this problem out please.

OpenStudy (unklerhaukus):

\[i^{12}=(i^2)^6=\dots\]

OpenStudy (anonymous):

\[\large i^2 = -1\]

OpenStudy (lgbasallote):

\[i^1 = i\] \[i^2 = -1\] \[i^3 = -i\] \[i^4 = 1\] does that help?

OpenStudy (anonymous):

Thank you. Could you explain if it would be different for i^49 also please?

OpenStudy (lgbasallote):

are you asking what the value of that is?

OpenStudy (lgbasallote):

just divide 49 by 4...then tell me the remainder

OpenStudy (anonymous):

It is Intermediate algebra. The directions just state to evaluate.

OpenStudy (anonymous):

12 remainder 1

OpenStudy (lgbasallote):

right so 1 is the remainder

OpenStudy (lgbasallote):

now look at the thing i posted earlier which corresponds to i^1

OpenStudy (unklerhaukus):

\[i^{49}=(i^4)^{12}\times i\]

OpenStudy (anonymous):

ok now that post makes sense thank you both.

OpenStudy (anonymous):

Now take out the remainder as the power of \(i\),, \[\large i^1 = ??\]

OpenStudy (anonymous):

i^49 = i correct

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

sorry -i

OpenStudy (anonymous):

How ??

OpenStudy (lgbasallote):

i^1 = i according to that post a while ago

OpenStudy (anonymous):

\[i*i ^{48}\] \[1*i=-i\]

OpenStudy (anonymous):

After finding the remainder just make the remainder as the power of \(i\) Here you got remainder as 1; \[\large i^1 = ??\]

OpenStudy (anonymous):

\[\huge 1 \times i = i \ne -i\]

OpenStudy (anonymous):

Not sure where I got the negative from. But your explanation makes more sense.

OpenStudy (anonymous):

So it is \(i\) and not \(-i\)..

OpenStudy (lgbasallote):

\[\Large i \times i^{48} \implies i \times (i^2)^{24} \implies i \times (-1)^{24} \implies i \times 1 \implies i\]

OpenStudy (lgbasallote):

does that make sense?

OpenStudy (anonymous):

Yes it does.